Product Cost Method of Product Pricing

La Femme Accessories Inc. produces women's handbags. The cost of producing 1,130 handbags is as follows:

Direct materials $17,000
Direct labor 7,100
Factory overhead 5,800
Total manufacturing cost $29,900

The selling and administrative expenses are $26,900. The management desires a profit equal to 15% of invested assets of $501,000.

If required, round your answers to nearest whole number.

a. Determine the amount of desired profit from the production and sale of 1,130 handbags.
$

b. Determine the product cost per unit for the production of 1,130 handbags.
$per unit

c. Determine the product cost markup percentage for handbags.
%

d. Determine the selling price of handbags. Round your answers to nearest whole value.

Cost $per unit
Markup $per unit
Selling price $per unit
Answer #1

solution a:

Amount of desired profit = $501,000*15% = $75,150

Solution b:

Product cost per unit = Total manufacturing costs / Nos of units produced = $29,900 / 1130 = $26.46 per unit

Solution c:

Product cost markup percentage = Required markup to cover selling and administrative expenses and profit / Total manufacturing costs

= ($26,900 + $75,150) / $29,900 = 341.30%

Solution d:

Selling price of handbags:

Cost = $26.46 per unit

Markup = $26.46 * 341.30% = $90.31 per unit

Selling price = $26.46 + $90.31 = $116.77 per unit = $117 per unit (rounded off)